# Relational Prompt-based Pre-trained Language Models for Social Event Detection

### v1

@@ 5.1. Model Effectiveness
$\mathrm{RPLM}_{SED}^{\mathrm{k}}$
$\mathrm{RPLM}_{SED}^{\mathrm{d}}$
Improve
$M_{1}$
.$23\pm.00$
.$36\pm.00$
.$39\pm.00$
.$43\pm.01$
.$38\pm.01$
.$\underline{.84\pm.01}$
.$46\pm.01$
.$\textbf{.91}\pm\textbf{.02}.$
.$\uparrow.07$
$M_{2}$
.$57\pm.00$
.$68\pm.00$
.$79\pm.01$
.$81\pm.02$
.$81\pm.00$
.$84\pm.01$
.$\underline{.89\pm.02}$
.$\textbf{.91}\pm\textbf{.01}.$
.$\uparrow.07$
$M_{3}$
.55±.01plus-or-minus.55.01.$5$01
.$63\pm.00$
.$67\pm.00$
.$78\pm.01$
.$71\pm.01$
.$\underline{.89\pm.00}$
.88±.00plus-or-minus.88.00.$8$00
.$\textbf{.93}\pm\textbf{.00}.$
.$\uparrow.04$
$M_{4}$
.$46\pm.01$
.$63\pm.00$
.$67\pm.02$
.$71\pm.02$
.$71\pm.01$
.$71\pm.01$
.$\underline{.78\pm.01}$
.$\textbf{.83}\pm\textbf{.01}.$
.$\uparrow.12$
$M_{5}$
.$48\pm.00$
.$59\pm.00$
.$73\pm.00$
.$75\pm.00$
.$76\pm.00$
.$83\pm.00$
.$\textbf{.85}\pm\textbf{.01}.$
.$\underline{.85\pm.02}$
.$\uparrow.02$
$M_{6}$
.$57\pm.01$
.$70\pm.00$
.$82\pm.01$
.$83\pm.01$
.$84\pm.00$
.$83\pm.00$
.$\underline{.87\pm.01}$
.$\textbf{.92}\pm\textbf{.00}.$
.$\uparrow.08$
$M_{7}$
.$37\pm.00$
.$51\pm.00$
.55±.01plus-or-minus.55.01.$5$01
.$57\pm.01$
.$56\pm.00$
.$\underline{.73\pm.01}$
.$68\pm.01$
[…]
.$\uparrow.15$
$M_{8}$
.55±.02plus-or-minus.55.02.$5$02
.$71\pm.00$
.$80\pm.00$
.$79\pm.01$
.$87\pm.01$
.$87\pm.02$
.$\underline{.88\pm.01}$
[…]
.$\uparrow.01$
$M_{9}$
.$51\pm.00$
.$67\pm.00$
.$74\pm.02$
.77±.02plus-or-minus.77.02.$7$02
.$78\pm.02$
.$79\pm.01$
.$\underline{.87\pm.01}$
.$\textbf{.92}\pm\textbf{.01}.$
.$\uparrow.13$
$M_{10}$
.55±.02plus-or-minus.55.02.$5$02
.$68\pm.00$
.$80\pm.01$
.$82\pm.02$
.$81\pm.01$
.$82\pm.01$
.$\underline{.90\pm.00}$
.$\textbf{.91}\pm\textbf{.01}.$
.$\uparrow.09$
$M_{11}$
.$50\pm.01$
.$65\pm.00$
.$74\pm.01$
.$75\pm.01$
.$76\pm.00$
.$75\pm.00$
.$\underline{.87\pm.01}$
[…]
.$\uparrow.12$
$M_{12}$
.$45\pm.01$
.$61\pm.00$
.$68\pm.01$
.$70\pm.01$
.$76\pm.01$
.$67\pm.01$
.$\underline{.85\pm.01}$
.$\textbf{.92}\pm\textbf{.00}.$
.$\uparrow.16$
$M_{13}$
.$47\pm.01$
.$58\pm.00$
.$69\pm.01$
.$68\pm.02$
.$67\pm.00$
.$\underline{.79\pm.00}$
.77±.01plus-or-minus.77.01.$7$01
.$\textbf{.91}\pm\textbf{.01}.$
.$\uparrow.12$
$M_{14}$
.44±.01plus-or-minus.44.01.$4$01
.$57\pm.00$
.$69\pm.00$
.$68\pm.01$
.$74\pm.00$
.$\underline{.82\pm.00}$
.$82\pm.01$
[…]
.$\uparrow.06$
$M_{15}$
.$39\pm.01$
.$49\pm.00$
.$58\pm.00$
.$59\pm.01$
.$64\pm.00$
.$69\pm.01$
.$\underline{.75\pm.01}$
.$\textbf{.83}\pm\textbf{.01}.$
.$\uparrow.14$
$M_{16}$
.55±.01plus-or-minus.55.01.$5$01
.$62\pm.00$
.$79\pm.01$
.$78\pm.01$
.$80\pm.00$
.$\underline{.90\pm.01}$
.$84\pm.00$
.$\textbf{.93}\pm\textbf{.01}.$
.$\uparrow.03$
$M_{17}$
.$48\pm.00$
.$58\pm.00$
.$70\pm.01$
.$71\pm.01$
.$73\pm.00$
.$\underline{.83\pm.00}$
.$80\pm.01$
.$\textbf{.86}\pm\textbf{.01}.$
.$\uparrow.03$
$M_{18}$
.$47\pm.01$
.$59\pm.00$
.$68\pm.02$
.$70\pm.01$
.$72\pm.01$
.$74\pm.01$
.$\underline{.82\pm.01}$
.$\textbf{.83}\pm\textbf{.00}.$
.$\uparrow.09$
$M_{19}$
.$51\pm.02$
.$60\pm.00$
.$73\pm.01$
.$73\pm.00$
.$76\pm.02$
.66±.01plus-or-minus.66.01.$6$01
.$\underline{.88\pm.01}$
.$\textbf{.91}\pm\textbf{.00}.$
.$\uparrow.15$
$M_{20}$
.$51\pm.01$
.$67\pm.00$
.$72\pm.02$
.$73\pm.02$
.$73\pm.00$
.$80\pm.00$
.$\textbf{.85}\pm\textbf{.00}.$
.$\underline{.82\pm.01}$
.$\uparrow.05$
$M_{21}$
.$41\pm.02$
.$53\pm.00$
.$60\pm.00$
.$61\pm.01$
.$65\pm.01$
.$\underline{.74\pm.01}$
.$\textbf{.76}\pm\textbf{.00}.$
.$74\pm.02$
.$\uparrow.02$
Blocks
PP-GCN
EventX
KPGNN
QSGNN
[…]
$\mathrm{RPLM}_{SED}^{\mathrm{k}}$
$\mathrm{RPLM}_{SED}^{\mathrm{d}}$
Improve
$M_{1}$
.$21\pm.00$
.$06\pm.00$
.$37\pm.00$
.$41\pm.02$
.$36\pm.01$
.$\underline{.84\pm.01}$
.44±.01plus-or-minus.44.01.$4$01
.$\textbf{.91}\pm\textbf{.01}.$
.$\uparrow.07$
$M_{2}$
.55±.02plus-or-minus.55.02.$5$02
.$29\pm.02$
.$78\pm.01$
.$80\pm.01$
.77±.00plus-or-minus.77.00.$7$00
.$84\pm.01$
.$\underline{.88\pm.01}$
.$\textbf{.91}\pm\textbf{.00}.$
.$\uparrow.07$
$M_{3}$
.$52\pm.01$
.$18\pm.01$
.$74\pm.00$
.$76\pm.01$
.$82\pm.01$
.$\underline{.89\pm.01}$
.$87\pm.00$
.$\textbf{.93}\pm\textbf{.00}.$
.$\uparrow.04$
$M_{4}$
.$42\pm.01$
.$19\pm.01$
.$64\pm.01$
.$68\pm.01$
.$67\pm.02$
.$69\pm.00$
.$\underline{.76\pm.01}$
.$\textbf{.81}\pm\textbf{.01}.$
.$\uparrow.12$
$M_{5}$
.$46\pm.01$
.$14\pm.00$
.$71\pm.01$
.$73\pm.00$
.$74\pm.00$
.$82\pm.00$
.$\underline{.83\pm.01}$
.$\textbf{.84}\pm\textbf{.02}.$
.$\uparrow.02$
$M_{6}$
.$52\pm.02$
.$27\pm.00$
.$79\pm.01$
.$80\pm.01$
.$81\pm.00$
.$82\pm.02$
.$\underline{.85\pm.01}$
.$\textbf{.91}\pm\textbf{.00}.$
.$\uparrow.09$
$M_{7}$
.$34\pm.00$
.$13\pm.00$
.$51\pm.01$
.$54\pm.00$
.$53\pm.00$
.$\underline{.72\pm.00}$
.66±.01plus-or-minus.66.01.$6$01
[…]
.$\uparrow.16$
$M_{8}$
.$49\pm.02$
.$21\pm.00$
.$76\pm.00$
.$75\pm.01$
.$84\pm.01$
.$\textbf{.87}\pm\textbf{.01}.$
.$86\pm.01$
.$\underline{.86\pm.00}$
.$\downarrow.01$
$M_{9}$
.$46\pm.02$
.$19\pm.00$
.$71\pm.02$
.$75\pm.02$
.$75\pm.00$
.$78\pm.01$
.$\underline{.86\pm.01}$
.$\textbf{.91}\pm\textbf{.00}.$
.$\uparrow.13$
$M_{10}$
.$51\pm.02$
.$24\pm.00$
.$78\pm.01$
.$80\pm.03$
.$78\pm.00$
.$81\pm.00$
.$\underline{.88\pm.01}$
.$\textbf{.90}\pm\textbf{.01}.$
.$\uparrow.09$
$M_{11}$
.$46\pm.01$
.$24\pm.00$
.$71\pm.01$
.$72\pm.01$
.$73\pm.00$
.$74\pm.00$
.$\underline{.85\pm.01}$
.$\textbf{.87}\pm\textbf{.01}.$
.$\uparrow.13$
$M_{12}$
.$42\pm.01$
.$16\pm.00$
.66±.01plus-or-minus.66.01.$6$01
.$68\pm.00$
.$75\pm.01$
.$67\pm.02$
.$\textbf{.84}\pm\textbf{.00}.$
.$\underline{.77\pm.00}$
.$\uparrow.09$
$M_{13}$
.$43\pm.01$
.$16\pm.00$
.$67\pm.01$
.66±.01plus-or-minus.66.01.$6$01
.$64\pm.00$
.$\underline{.79\pm.00}$
.$75\pm.01$
.$\textbf{.91}\pm\textbf{.00}.$
.$\uparrow.12$
$M_{14}$
.$41\pm.01$
.$14\pm.00$
.$65\pm.00$
.66±.01plus-or-minus.66.01.$6$01
.$72\pm.00$
.$\underline{.82\pm.01}$
.$81\pm.01$
[…]
.$\uparrow.06$
$M_{15}$
.$35\pm.01$
.$07\pm.00$
.$54\pm.00$
.55±.01plus-or-minus.55.01.$5$01
.$61\pm.00$
.$67\pm.01$
.$\underline{.73\pm.01}$
.$\textbf{.82}\pm\textbf{.01}.$
.$\uparrow.15$
$M_{16}$
.$52\pm.01$
.$19\pm.00$
.77±.01plus-or-minus.77.01.$7$01
.$76\pm.02$
.$75\pm.01$
.$\underline{.90\pm.00}$
.$82\pm.00$
.$\textbf{.93}\pm\textbf{.01}.$
.$\uparrow.03$
$M_{17}$
.$45\pm.00$
.$18\pm.00$
.$68\pm.01$
.$69\pm.01$
.$71\pm.02$
.$\underline{.82\pm.00}$
.$79\pm.00$
.$\textbf{.86}\pm\textbf{.01}.$
.$\uparrow.04$
$M_{18}$
.$45\pm.01$
.$16\pm.00$
.66±.02plus-or-minus.66.02.$6$02
.$68\pm.01$
.$70\pm.00$
.$74\pm.00$
.$\underline{.81\pm.01}$
.$\textbf{.82}\pm\textbf{.01}.$
.$\uparrow.08$
$M_{19}$
.$48\pm.02$
.$16\pm.00$
.$71\pm.01$
.$70\pm.01$
.$75\pm.01$
.66±.00plus-or-minus.66.00.$6$00
.$\underline{.86\pm.01}$
.$\textbf{.90}\pm\textbf{.01}.$
.$\uparrow.15$
$M_{20}$
.$45\pm.02$
.$18\pm.00$
.$68\pm.02$
.$69\pm.02$
.$68\pm.00$
.$78\pm.00$
.$\textbf{.82}\pm\textbf{.00}.$
.$\underline{.80\pm.01}$
.$\uparrow.04$
$M_{21}$
.$38\pm.02$
.$10\pm.00$
.$57\pm.00$
.$58\pm.00$
.$63\pm.01$
.$64\pm.01$
.$\textbf{.74}\pm\textbf{.01}.$
.$\underline{.71\pm.01}$
.$\uparrow.10$
Blocks
PP-GCN
EventX
KPGNN
QSGNN
[…]
$\mathrm{RPLM}_{SED}^{\mathrm{k}}$
$\mathrm{RPLM}_{SED}^{\mathrm{d}}$
Improve
$M_{1}$
.$05\pm.00$
.$01\pm.00$
.$07\pm.01$
.$09\pm.01$
.$05\pm.00$
.$\underline{.90\pm.00}$
.$07\pm.01$
.$\textbf{.96}\pm\textbf{.01}.$
.$\uparrow.06$
$M_{2}$
.$67\pm.03$
.$45\pm.02$
.$76\pm.02$
.$78\pm.02$
.$67\pm.01$
.$90\pm.01$
.$83\pm.02$
.$\textbf{.91}\pm\textbf{.01}.$
.$\uparrow.00$
$M_{3}$
.$47\pm.01$
.$09\pm.01$
.$58\pm.01$
.$62\pm.01$
.$58\pm.00$
.$\underline{.89\pm.01}$
.66±.01plus-or-minus.66.01.$6$01
.$\textbf{.94}\pm\textbf{.00}.$
.$\uparrow.05$
$M_{4}$
.$24\pm.01$
.$07\pm.01$
.$29\pm.01$
.$31\pm.02$
.$27\pm.02$
.$27\pm.01$
.$\underline{.39\pm.02}$
.$\textbf{.61}\pm\textbf{.00}.$
.$\uparrow.32$
$M_{5}$
.$34\pm.00$
.$04\pm.00$
.$47\pm.03$
.$48\pm.02$
.$43\pm.01$
.$63\pm.02$
.$\underline{.67\pm.01}$
.$\textbf{.74}\pm\textbf{.02}.$
.$\uparrow.11$
$M_{6}$
.55±.03plus-or-minus.55.03.$5$03
.$14\pm.00$
.$72\pm.03$
.$65\pm.03$
.$65\pm.00$
.$\underline{.74\pm.00}$
.$65\pm.01$
.$\textbf{.93}\pm\textbf{.00}.$
.$\uparrow.19$
$M_{7}$
.11±.02plus-or-minus.11.02.$1$02
.$02\pm.00$
.$12\pm.00$
.$13\pm.02$
.$09\pm.01$
.$\underline{.45\pm.01}$
.$19\pm.01$
.$\textbf{.91}\pm\textbf{.01}.$
.$\uparrow.46$
$M_{8}$
.$43\pm.04$
.$09\pm.00$
.$60\pm.01$
.$57\pm.02$
.$65\pm.02$
.$\underline{.72\pm.01}$
.$68\pm.02$
.$\textbf{.79}\pm\textbf{.01}.$
.$\uparrow.07$
$M_{9}$
.$31\pm.02$
.$07\pm.00$
.$46\pm.02$
.$48\pm.03$
.$43\pm.00$
.$\underline{.68\pm.00}$
.$56\pm.01$
.$\textbf{.95}\pm\textbf{.01}.$
.$\uparrow.27$
$M_{10}$
.$50\pm.07$
.$13\pm.00$
.$70\pm.06$
.$65\pm.02$
.$62\pm.02$
.$74\pm.00$
.$\underline{.84\pm.01}$
.$\textbf{.93}\pm\textbf{.01}.$
.$\uparrow.19$
$M_{11}$
.$38\pm.02$
.$16\pm.00$
.$49\pm.03$
.$50\pm.03$
.$42\pm.01$
.$60\pm.00$
.$\underline{.61\pm.06}$
.$\textbf{.91}\pm\textbf{.01}.$
.$\uparrow.31$
$M_{12}$
.$34\pm.03$
.$07\pm.00$
.$48\pm.01$
.$47\pm.01$
.44±.00plus-or-minus.44.00.$4$00
.$26\pm.00$
.$\textbf{.69}\pm\textbf{.03}.$
.$\underline{.48\pm.03}$
.$\uparrow.21$
$M_{13}$
.$19\pm.01$
.$04\pm.00$
.$29\pm.03$
.$28\pm.02$
.$21\pm.02$
.$\underline{.75\pm.02}$
.$32\pm.03$
.$\textbf{.97}\pm\textbf{.00}.$
.$\uparrow.22$
$M_{14}$
.$29\pm.01$
.$10\pm.00$
.$42\pm.02$
$39\pm.01$
.$43\pm.01$
.$\underline{.81\pm.01}$
.$53\pm.02$
.$\textbf{.85}\pm\textbf{.01}.$
.$\uparrow.04$
$M_{15}$
.$15\pm.00$
.$01\pm.00$
.$17\pm.00$
.$15\pm.01$
.$16\pm.00$
.$\underline{.46\pm.00}$
.$31\pm.02$
.$\textbf{.78}\pm\textbf{.00}.$
.$\uparrow.32$
$M_{16}$
.$51\pm.03$
.$08\pm.00$
.66±.05plus-or-minus.66.05.$6$05
.$58\pm.04$
.$56\pm.01$
.$\underline{.88\pm.01}$
.$61\pm.01$
.$\textbf{.95}\pm\textbf{.01}.$
.$\uparrow.07$
$M_{17}$
.$35\pm.03$
.$12\pm.00$
.$43\pm.05$
.$40\pm.02$
.$36\pm.01$
.$\underline{.81\pm.01}$
.44±.01plus-or-minus.44.01.$4$01
.$\textbf{.89}\pm\textbf{.00}.$
.$\uparrow.08$
$M_{18}$
.$39\pm.03$
.$08\pm.00$
.$47\pm.04$
.$48\pm.01$
.44±.01plus-or-minus.44.01.$4$01
.$52\pm.01$
.$\underline{.56\pm.01}$
.$\textbf{.67}\pm\textbf{.02}.$
.$\uparrow.15$
$M_{19}$
.$41\pm.02$
.$07\pm.00$
.$51\pm.03$
.$45\pm.03$
.44±.00plus-or-minus.44.00.$4$00
.$35\pm.01$
.$\underline{.62\pm.03}$
.$\textbf{.90}\pm\textbf{.01}.$
.$\uparrow.39$
$M_{20}$
.$41\pm.01$
.11±.00plus-or-minus.11.00.$1$00
.$51\pm.04$
.$47\pm.02$
.$43\pm.02$
.$\underline{.71\pm.01}$
.$64\pm.02$
.$\textbf{.75}\pm\textbf{.02}.$
.$\uparrow.04$
$M_{21}$
.$20\pm.03$
.$01\pm.00$
.$20\pm.01$
.$24\pm.01$
.$23\pm.00$
.$\textbf{.48}\pm\textbf{.00}.$
.$37\pm.02$
.$\underline{.44\pm.01}$
.$\downarrow.04$

### v2

@@ 5.1. Model Effectiveness
$M_{1}$$23\pm.00$$36\pm.00$$39\pm.00$$43\pm.01$$38\pm.01$$\underline{.84\pm.01}$$46\pm.01$$\textbf{.91}\pm\textbf{.02}.$$\uparrow.07$
$M_{2}$$57\pm.00$$68\pm.00$$79\pm.01$$81\pm.02$$81\pm.00$$84\pm.01$$\underline{.89\pm.02}$$\textbf{.91}\pm\textbf{.01}.$$\uparrow.07$
$M_{3}$$5$01 .$63\pm.00$$67\pm.00$$78\pm.01$$71\pm.01$$\underline{.89\pm.00}$$8$00 .$\textbf{.93}\pm\textbf{.00}.$$\uparrow.04$
$M_{4}$$46\pm.01$$63\pm.00$$67\pm.02$$71\pm.02$$71\pm.01$$71\pm.01$$\underline{.78\pm.01}$$\textbf{.83}\pm\textbf{.01}.$$\uparrow.12$
$M_{5}$$48\pm.00$$59\pm.00$$73\pm.00$$75\pm.00$$76\pm.00$$83\pm.00$$\textbf{.85}\pm\textbf{.01}.$$\underline{.85\pm.02}$$\uparrow.02$
$M_{6}$$57\pm.01$$70\pm.00$$82\pm.01$$83\pm.01$$84\pm.00$$83\pm.00$$\underline{.87\pm.01}$$\textbf{.92}\pm\textbf{.00}.$$\uparrow.08$
$M_{7}$$37\pm.00$$51\pm.00$$5$01 .$57\pm.01$$56\pm.00$$\underline{.73\pm.01}$$68\pm.01$$8$01 .$\uparrow.15$
$M_{8}$$5$02 .$71\pm.00$$80\pm.00$$79\pm.01$$87\pm.01$$87\pm.02$$\underline{.88\pm.01}$$8$00 .$\uparrow.01$
$M_{9}$$51\pm.00$$67\pm.00$$74\pm.02$$7$02 .$78\pm.02$$79\pm.01$$\underline{.87\pm.01}$$\textbf{.92}\pm\textbf{.01}.$$\uparrow.13$
$M_{10}$$5$02 .$68\pm.00$$80\pm.01$$82\pm.02$$81\pm.01$$82\pm.01$$\underline{.90\pm.00}$$\textbf{.91}\pm\textbf{.01}.$$\uparrow.09$
$M_{11}$$50\pm.01$$65\pm.00$$74\pm.01$$75\pm.01$$76\pm.00$$75\pm.00$$\underline{.87\pm.01}$$8$01 .$\uparrow.12$
$M_{12}$$45\pm.01$$61\pm.00$$68\pm.01$$70\pm.01$$76\pm.01$$67\pm.01$$\underline{.85\pm.01}$$\textbf{.92}\pm\textbf{.00}.$$\uparrow.16$
$M_{13}$$47\pm.01$$58\pm.00$$69\pm.01$$68\pm.02$$67\pm.00$$\underline{.79\pm.00}$$7$01 .$\textbf{.91}\pm\textbf{.01}.$$\uparrow.12$
$M_{14}$$4$01 .$57\pm.00$$69\pm.00$$68\pm.01$$74\pm.00$$\underline{.82\pm.00}$$82\pm.01$$8$00 .$\uparrow.06$
$M_{15}$$39\pm.01$$49\pm.00$$58\pm.00$$59\pm.01$$64\pm.00$$69\pm.01$$\underline{.75\pm.01}$$\textbf{.83}\pm\textbf{.01}.$$\uparrow.14$
$M_{16}$$5$01 .$62\pm.00$$79\pm.01$$78\pm.01$$80\pm.00$$\underline{.90\pm.01}$$84\pm.00$$\textbf{.93}\pm\textbf{.01}.$$\uparrow.03$
$M_{17}$$48\pm.00$$58\pm.00$$70\pm.01$$71\pm.01$$73\pm.00$$\underline{.83\pm.00}$$80\pm.01$$\textbf{.86}\pm\textbf{.01}.$$\uparrow.03$
$M_{18}$$47\pm.01$$59\pm.00$$68\pm.02$$70\pm.01$$72\pm.01$$74\pm.01$$\underline{.82\pm.01}$$\textbf{.83}\pm\textbf{.00}.$$\uparrow.09$
$M_{19}$$51\pm.02$$60\pm.00$$73\pm.01$$73\pm.00$$76\pm.02$$6$01 .$\underline{.88\pm.01}$$\textbf{.91}\pm\textbf{.00}.$$\uparrow.15$
$M_{20}$$51\pm.01$$67\pm.00$$72\pm.02$$73\pm.02$$73\pm.00$$80\pm.00$$\textbf{.85}\pm\textbf{.00}.$$\underline{.82\pm.01}$$\uparrow.05$
$M_{21}$$41\pm.02$$53\pm.00$$60\pm.00$$61\pm.01$$65\pm.01$$\underline{.74\pm.01}$$\textbf{.76}\pm\textbf{.00}.$$74\pm.02$$\uparrow.02$
@@ 5.1.1. Offline Evaluation
[…]
$M_{1}$$21\pm.00$$06\pm.00$$37\pm.00$$41\pm.02$$36\pm.01$$\underline{.84\pm.01}$$4$01 .$\textbf{.91}\pm\textbf{.01}.$$\uparrow.07$
$M_{2}$$5$02 .$29\pm.02$$78\pm.01$$80\pm.01$$7$00 .$84\pm.01$$\underline{.88\pm.01}$$\textbf{.91}\pm\textbf{.00}.$$\uparrow.07$
$M_{3}$$52\pm.01$$18\pm.01$$74\pm.00$$76\pm.01$$82\pm.01$$\underline{.89\pm.01}$$87\pm.00$$\textbf{.93}\pm\textbf{.00}.$$\uparrow.04$
$M_{4}$$42\pm.01$$19\pm.01$$64\pm.01$$68\pm.01$$67\pm.02$$69\pm.00$$\underline{.76\pm.01}$$\textbf{.81}\pm\textbf{.01}.$$\uparrow.12$
$M_{5}$$46\pm.01$$14\pm.00$$71\pm.01$$73\pm.00$$74\pm.00$$82\pm.00$$\underline{.83\pm.01}$$\textbf{.84}\pm\textbf{.02}.$$\uparrow.02$
$M_{6}$$52\pm.02$$27\pm.00$$79\pm.01$$80\pm.01$$81\pm.00$$82\pm.02$$\underline{.85\pm.01}$$\textbf{.91}\pm\textbf{.00}.$$\uparrow.09$
[…]
$M_{8}$$49\pm.02$$21\pm.00$$76\pm.00$$75\pm.01$$84\pm.01$$\textbf{.87}\pm\textbf{.01}.$$86\pm.01$$\underline{.86\pm.00}$$\downarrow.01$
$M_{9}$$46\pm.02$$19\pm.00$$71\pm.02$$75\pm.02$$75\pm.00$$78\pm.01$$\underline{.86\pm.01}$$\textbf{.91}\pm\textbf{.00}.$$\uparrow.13$
$M_{10}$$51\pm.02$$24\pm.00$$78\pm.01$$80\pm.03$$78\pm.00$$81\pm.00$$\underline{.88\pm.01}$$\textbf{.90}\pm\textbf{.01}.$$\uparrow.09$
$M_{11}$$46\pm.01$$24\pm.00$$71\pm.01$$72\pm.01$$73\pm.00$$74\pm.00$$\underline{.85\pm.01}$$\textbf{.87}\pm\textbf{.01}.$$\uparrow.13$
$M_{12}$$42\pm.01$$16\pm.00$$6$01 .$68\pm.00$$75\pm.01$$67\pm.02$$\textbf{.84}\pm\textbf{.00}.$$\underline{.77\pm.00}$$\uparrow.09$
$M_{13}$$43\pm.01$$16\pm.00$$67\pm.01$$6$01 .$64\pm.00$$\underline{.79\pm.00}$$75\pm.01$$\textbf{.91}\pm\textbf{.00}.$$\uparrow.12$
$M_{14}$$41\pm.01$$14\pm.00$$65\pm.00$$6$01 .$72\pm.00$$\underline{.82\pm.01}$$81\pm.01$$8$01 .$\uparrow.06$
$M_{15}$$35\pm.01$$07\pm.00$$54\pm.00$$5$01 .$61\pm.00$$67\pm.01$$\underline{.73\pm.01}$$\textbf{.82}\pm\textbf{.01}.$$\uparrow.15$
$M_{16}$$52\pm.01$$19\pm.00$$7$01 .$76\pm.02$$75\pm.01$$\underline{.90\pm.00}$$82\pm.00$$\textbf{.93}\pm\textbf{.01}.$$\uparrow.03$
$M_{17}$$45\pm.00$$18\pm.00$$68\pm.01$$69\pm.01$$71\pm.02$$\underline{.82\pm.00}$$79\pm.00$$\textbf{.86}\pm\textbf{.01}.$$\uparrow.04$
$M_{18}$$45\pm.01$$16\pm.00$$6$02 .$68\pm.01$$70\pm.00$$74\pm.00$$\underline{.81\pm.01}$$\textbf{.82}\pm\textbf{.01}.$$\uparrow.08$
$M_{19}$$48\pm.02$$16\pm.00$$71\pm.01$$70\pm.01$$75\pm.01$$6$00 .$\underline{.86\pm.01}$$\textbf{.90}\pm\textbf{.01}.$$\uparrow.15$
$M_{20}$$45\pm.02$$18\pm.00$$68\pm.02$$69\pm.02$$68\pm.00$$78\pm.00$$\textbf{.82}\pm\textbf{.00}.$$\underline{.80\pm.01}$$\uparrow.04$
$M_{21}$$38\pm.02$$10\pm.00$$57\pm.00$$58\pm.00$$63\pm.01$$64\pm.01$$\textbf{.74}\pm\textbf{.01}.$$\underline{.71\pm.01}$$\uparrow.10$
[…]
$M_{1}$$05\pm.00$$01\pm.00$$07\pm.01$$09\pm.01$$05\pm.00$$\underline{.90\pm.00}$$07\pm.01$$\textbf{.96}\pm\textbf{.01}.$$\uparrow.06$
$M_{2}$$67\pm.03$$45\pm.02$$76\pm.02$$78\pm.02$$67\pm.01$$90\pm.01$$83\pm.02$$\textbf{.91}\pm\textbf{.01}.$$\uparrow.00$
$M_{3}$$47\pm.01$$09\pm.01$$58\pm.01$$62\pm.01$$58\pm.00$$\underline{.89\pm.01}$$6$01 .$\textbf{.94}\pm\textbf{.00}.$$\uparrow.05$
$M_{4}$$24\pm.01$$07\pm.01$$29\pm.01$$31\pm.02$$27\pm.02$$27\pm.01$$\underline{.39\pm.02}$$\textbf{.61}\pm\textbf{.00}.$$\uparrow.32$
$M_{5}$$34\pm.00$$04\pm.00$$47\pm.03$$48\pm.02$$43\pm.01$$63\pm.02$$\underline{.67\pm.01}$$\textbf{.74}\pm\textbf{.02}.$$\uparrow.11$
$M_{6}$$5$03 .$14\pm.00$$72\pm.03$$65\pm.03$$65\pm.00$$\underline{.74\pm.00}$$65\pm.01$$\textbf{.93}\pm\textbf{.00}.$$\uparrow.19$
$M_{7}$$1$02 .$02\pm.00$$12\pm.00$$13\pm.02$$09\pm.01$$\underline{.45\pm.01}$$19\pm.01$$\textbf{.91}\pm\textbf{.01}.$$\uparrow.46$
$M_{8}$$43\pm.04$$09\pm.00$$60\pm.01$$57\pm.02$$65\pm.02$$\underline{.72\pm.01}$$68\pm.02$$\textbf{.79}\pm\textbf{.01}.$$\uparrow.07$
$M_{9}$$31\pm.02$$07\pm.00$$46\pm.02$$48\pm.03$$43\pm.00$$\underline{.68\pm.00}$$56\pm.01$$\textbf{.95}\pm\textbf{.01}.$$\uparrow.27$
$M_{10}$$50\pm.07$$13\pm.00$$70\pm.06$$65\pm.02$$62\pm.02$$74\pm.00$$\underline{.84\pm.01}$$\textbf{.93}\pm\textbf{.01}.$$\uparrow.19$
$M_{11}$$38\pm.02$$16\pm.00$$49\pm.03$$50\pm.03$$42\pm.01$$60\pm.00$$\underline{.61\pm.06}$$\textbf{.91}\pm\textbf{.01}.$$\uparrow.31$
$M_{12}$$34\pm.03$$07\pm.00$$48\pm.01$$47\pm.01$$4$00 .$26\pm.00$$\textbf{.69}\pm\textbf{.03}.$$\underline{.48\pm.03}$$\uparrow.21$
$M_{13}$$19\pm.01$$04\pm.00$$29\pm.03$$28\pm.02$$21\pm.02$$\underline{.75\pm.02}$$32\pm.03$$\textbf{.97}\pm\textbf{.00}.$$\uparrow.22$
$M_{14}$$29\pm.01$$10\pm.00$$42\pm.02$ $39\pm.01$$43\pm.01$$\underline{.81\pm.01}$$53\pm.02$$\textbf{.85}\pm\textbf{.01}.$$\uparrow.04$
$M_{15}$$15\pm.00$$01\pm.00$$17\pm.00$$15\pm.01$$16\pm.00$$\underline{.46\pm.00}$$31\pm.02$$\textbf{.78}\pm\textbf{.00}.$$\uparrow.32$
$M_{16}$$51\pm.03$$08\pm.00$$6$05 .$58\pm.04$$56\pm.01$$\underline{.88\pm.01}$$61\pm.01$$\textbf{.95}\pm\textbf{.01}.$$\uparrow.07$
$M_{17}$$35\pm.03$$12\pm.00$$43\pm.05$$40\pm.02$$36\pm.01$$\underline{.81\pm.01}$$4$01 .$\textbf{.89}\pm\textbf{.00}.$$\uparrow.08$
$M_{18}$$39\pm.03$$08\pm.00$$47\pm.04$$48\pm.01$$4$01 .$52\pm.01$$\underline{.56\pm.01}$$\textbf{.67}\pm\textbf{.02}.$$\uparrow.15$
$M_{19}$$41\pm.02$$07\pm.00$$51\pm.03$$45\pm.03$$4$00 .$35\pm.01$$\underline{.62\pm.03}$$\textbf{.90}\pm\textbf{.01}.$$\uparrow.39$
$M_{20}$$41\pm.01$$1$00 .$51\pm.04$$47\pm.02$$43\pm.02$$\underline{.71\pm.01}$$64\pm.02$$\textbf{.75}\pm\textbf{.02}.$$\uparrow.04$
$M_{21}$$20\pm.03$$01\pm.00$$20\pm.01$$24\pm.01$$23\pm.00$$\textbf{.48}\pm\textbf{.00}.$$37\pm.02$$\underline{.44\pm.01}$$\downarrow.04$